1) What two sources of error are most likely to affect your results?
2) Suppose protons were emitted rather than electrons. How would this affect the experiment?
3) Compute the velocity of an electron that has been accelerated through a difference of
potential of 100 volts. Express your answer in meters per second.
4) Find the mass of the electron from your e/m and the accepted value of the electron’s charge,
e=1.602x10 ^-19 C
Answer #1

Please Specify which experiment you are talking about.

Question 3 can be answered however.

velocity of the electron is given by

v = [2eV/m]1/2

e = magnitude of electron charge = 1.6*10-19

m = mass of the electron = 9.1*10-31

V = accelerating voltage = 100 volts

v = 1.874 * 107 m/s